# Erasing Conceptual Knowledge from Language Models

### v2

@@ Abstract
In this work, we propose Erasure of Language Memory (ELM), an approach to concept-level unlearning built on the principle of matching the distribution defined by an introspective classifier. Our key insight is that effective unlearning should leverage the model’s ability to evaluate its own knowledge, using the model itself as a classifier to identify and reduce the likelihood of generating content related to undesired concepts. ELM applies this framework to create targeted low-rank updates that reduce generation probabilities for concept-specific content while preserving the model’s broader capabilities. We demonstrate ELM’s efficacy on biosecurity, cybersecurity, and literary domain erasure tasks. Comparative analysis shows that ELM achieves superior performance across key metrics, including near-random scores on erased topic assessments, maintained coherence in text generation, preserved accuracy on unrelated benchmarks, and robustness under adversarial attacks. Our code, data, and trained models are available at elm.baulab.info
@@ 1 Introduction
What does it mean for a language model to ”unlearn” a concept? While machine unlearning has traditionally focused on removing specific training samples from model memory, there is an increasing need to be able to erase broad conceptual knowledge—for example, removing all information about biological weapons rather than just a few training examples containing that information. In this paper we examine how concept-level unlearning leads to a new approach to knowledge removal in language models.
Prior approaches to unlearning broadly fall into three categories: (1) retraining on filtered data (2) reversed-gradient-based methods that attempt to ”un-train” specific knowledge, and (3) representation manipulation approaches that disrupt internal activations for targeted content. Unfortunately, each of these strategies have limitations that make them impractical for unlearning in large language models: dataset filtering requires retraining that is costly at scale; gradient reversal methods are unstable and create broad damage to the model; and representation manipulation creates obvious behavioral artifacts. These approaches lack a principled objective defining successful concept erasure. They focus on technical mechanisms like reversing gradients, altering training data, or randomizing activations without a clear target for the model’s modified behavior.
We propose a fundamentally different approach that leverages the model’s own ability to recognize and classify knowledge. Our key insight is that language models can act as their own critics: they can evaluate whether a piece of text demonstrates knowledge of a particular concept. This self-classification provides a natural objective for unlearning: we can modify the model to reduce the likelihood of generating text it would classify as containing target concept.
This insight leads to Erasure of Language Memory (ELM), a method that directly optimizes the model’s generation probabilities based on introspective classification. Unlike approaches like Representation Misdirection for Unlearning (RMU; Li et al., 2024) which manipulates internal activations without a clear behavioral target, or WhoIsHarryPotter (Eldan & Russinovich, 2023) which modifies training data but fails to fully eliminate concept knowledge, ELM has a principled objective: the model should generate coherent text that the language model itself would not classify as demonstrating knowledge of the target concept.
@@
Prior unlearning methods face fundamental limitations when evaluated holistically. Our proposed evaluation desiderata—innocence (lack of target knowledge), specificity (preserved capabilities), and seamlessness (coherent generation)—reveal these tradeoffs: gradient reversal achieves innocence but degrades general capabilities, representation manipulation preserves capabilities but generates incoherent text, and dataset filtering maintains coherence but fails to fully eliminate knowledge.
Through extensive experiments on WMDP biosecurity and cybersecurity benchmarks, as well as literary domain erasure, we measure the effectiveness of ELM’s principled approach to concept removal compared to previous methods. Our evaluation using both standard metrics and adversarial probing shows that ELM achieves robust concept erasure while maintaining model coherence and general capabilities.
@@ 2 Related work
The idea of removing specific data from machine learning models, known as machine unlearning, has gained attention in recent years, initially motivated by privacy concerns (Cao & Yang, 2015; Harding et al., 2019). Early methods focused on efficiently removing individual training examples or facts from models (Golatkar et al., 2020; Ma et al., 2022; Jang et al., 2022a). However, most existing benchmarks evaluate unlearning on artificially created deletion sets (Choi & Na, 2023; Goel et al., 2022; Maini et al., 2024), in contrast to our focus on real-world distributions of broad conceptual knowledge.
@@
New approaches to machine unlearning have recently gained traction on the problem of removing dangerous capabilities from LLMs (Lynch et al., 2024; Ilharco et al., 2023; Jang et al., 2022b; Lu et al., 2022; Yu et al., 2023; Casper et al., 2024; Eldan & Russinovich, 2023). Our work is most directly comparable to three state-of-the-art techniques: Representation Misdirection for Unlearning (RMU) (Li et al., 2024), WhoIsHarryPotter (WHP) (Eldan & Russinovich, 2023), and Representation Noising (RepNoise) (Rosati et al., 2024). RMU fine-tunes models to align internal activations with random scaled vectors when processing targeted concepts. WHP (Eldan & Russinovich, 2023) employs a two-stage approach, first training a reinforced model for the concept being erased and then training an unlearned model to behave differently on the reinforced logits. RepNoise (Rosati et al., 2024) focuses on removing information about harmful representations across all layers of the LLM, by doing gradient ascent along with some representation noising (training internal representations to match Gaussian noise). While all these methods successfully reduce model performance on erased knowledge, our measurements reveal that these previous approaches fall short in one or more of the three erasing goals. In contrast, our current work erases concepts by fine-tuning towards a principled target distribution that is designed to balance all three desiderata of innocence, specificity, and seamlessness.
Other methods such as LLMU (Yao et al., 2023), SSD (Foster et al., 2024), and SCRUB (Kurmanji et al., 2024) have also been proposed but face significant limitations. LLMU uses a differentiable membership classifier to identify unlearning set examples but struggles with imprecisely defined target distributions (see Li et al., 2024). SSD offers a retraining-free method for removing specific samples but is limited to individual examples rather than broader knowledge domains. SCRUB aims to remove sensitive data while preserving model performance on other tasks but requires access to the full training dataset, which is often impractical. Importantly, these methods were compared by RMU (Li et al., 2024) and found to be less effective for erasing broad conceptual knowledge.
@@
Some techniques for controlling generative model outputs involve using an auxiliary generative model to specify the desired behavior over some training data, and then distilling these outputs into a target model by training the target model to mimic the demonstrated behavior. For example Askell et al. (2021) and Bai et al. (2022) prompt an unsafe model into behaving more safely, and then distil the prompted model. Similarly, Gandikota et al. (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes. (Rosati et al., 2024) can also be classified under distillation methods as it mimics a Gaussian distribution when processing harmful tokens. Similarly, ELM matches the harmful logits to a modified output distribution, but unlike the distillation methods that usually adhere to the standard matching objective, we employ a multi-objective framework to address the seamlessness and specificity concerns with standard distillation methods. Prior works like Emulator and DeRa leverage probability ratios for behavioral modification - Emulator (Mitchell et al., 2023) scales fine-tuning changes between model sizes, while DeRa (Liu et al., 2024) controls alignment strength during inference. In contrast, ELM introduces a simpler, principled approach focused specifically on reducing the likelihood of knowledge concept generation.
@@
Gandikota et al. (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes. Gandikota et al. (2024) erase concepts by modifying the key value mapping of cross attention layers in a low rank closed form update. Other works remove the knowledge of unwanted concepts from the model weights; proposing attention re-steering through fine-tuning (Zhang et al., 2023), fine-tuning the attention weights (Kumari et al., 2023) and continual learning (Heng & Soh, 2023). We take inspiration from (Gandikota et al., 2023) to reduce the likelihood of a concept being generated.
3 Background
3.1 Transformer
We focus on autoregressive transformer-based models (Tunstall et al., 2023a; Dubey et al., 2024). Given an input token sequence , these models output a probability distribution over a vocabulary to predict the next token:
The model architecture comprises stacked blocks, each usually containing a multi-head attention (MHA) layer followed by a multi-layer perceptron (MLP). These layers are stacked on top of each other, and their outputs are added to the residual stream of the model. MHA layers play an important role in bringing information from previous input tokens, while MLP layers are found to play the role of knowledge banks (Meng et al., 2022; Geva et al., 2023). These calculations can be summarized as follows,
where denotes the hidden state for token at layer .
Previous research (Meng et al., 2022) has localized much model knowledge within early to mid-layer blocks. As our own extensive experiments have similarly found good efficacy when erasing from earlier blocks of parameters, we restrict our parameter changes to MLP and MHA layers from those same blocks.
3.2 Transformer Next Token Prediction: A Classification Perspective
Language models are typically viewed through the lens of autoregressive sequence modeling, where the model predicts the next token given a sequence of previous tokens. However, these models can also be understood as powerful classifiers. To illustrate this connection, let’s examine both perspectives:
Traditional Perspective: The standard way to describe an autoregressive language model is through the factorization:
@@ 3.2 Transformer Next Token Prediction: A Classification Perspective
Classification Perspective: We can reformulate this using Bayes’ rule to obtain:
In this view, the model acts as a classifier where previous tokens serve as class labels for the text that follows. For instance, if we provide the context as ”The following text is about general biology”, the term represents the probability that the generated text belongs to the ”general biology” class.
This classification perspective enables us to manipulate the model’s knowledge by adjusting these classification probabilities. We can define a target distribution that modifies how strongly the model associates generated text with specific concepts:
where is a scaling factor that controls the strength of the modification. When , we increase the likelihood of generating text associated with the class ; when , we decrease it.
To implement this in practice with an autoregressive language model, we apply Bayes’ rule again:
In this form, both the numerator and denominator can be read out of the pretrained language model. Eq 6 is particularly powerful for concept unlearning because it allows us to use the language model itself as both the generator and classifier of content. For example, to unlearn bioweapons expertise, we can reduce the model’s probability of generating text that would naturally follow the context ”You are an expert in bioweapons.” This approach requires no external classifier - we leverage the model’s own understanding of concept relationships as encoded in its next-token predictions. Our perspective is inspired by the classifier-free guidance work (Ho & Salimans, 2022) in diffusion model research which has later been adopted to language models (Sanchez et al., 2023).
3.3 Parameter-Efficient Fine-tuning via Low-Rank Adapters
Fine-tuning large pre-trained language models involves large sets of parameters, which can result in overfitting. To address this issue, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021), a method that introduces a factorization of weights to significantly reduce the number of parameters. Consider a pre-trained model layer with weights , where and denote input and output dimensions, respectively. LoRA decomposes the weight update as follows:
Here, represents a small rank (), constraining the update to a low-dimensional subspace. By optimizing only and while keeping fixed, LoRA significantly reduces the trainable parameters. During inference, the update can be efficiently incorporated into the weights:
The LoRA approach offers a balance between adaptation capability, reduced overfitting, and parameter efficiency, making it particularly suitable for our setting of fine-tuning large-scale models.
@@ 4 Method
We introduce Erasure of Language Memory (ELM), an approach that reformulates concept unlearning through introspective classification. While traditional unlearning methods have focused on sample removal through dataset retraining, gradient ascent, or representation disruption, ELM leverages the language model’s own ability to evaluate and modify its knowledge.
Specifically, ELM uses the language model itself as its own introspective classifier, as established in Section 3.2. Rather than training external classifiers or explicitly manipulating representations, we leverage the model’s next-token predictions to identify and reduce the likelihood of generating concept-specific content. This allows us to operate directly on the model’s probability distributions, providing a principled approach to unlearning.
@@
The core of our method is a self-classification objective that reduces the likelihood of generating text that the model would classify as containing the target concept. Given a dataset containing text related to the target concept, we modify the model’s predicted probabilities to diverge from generating content related to the concept being erased and instead favor generating content related to a safer alternative :
where controls the strength of knowledge modification. is the concept to be erased, and is an alternative concept. The corresponding loss is:
@@
In practice, we encounter two main challenges when implementing this objective. First, knowledge in language models is often entangled - modifying one concept can unintentionally affect related concepts. To address this, we preserve the model’s behavior on a set of related but safe concepts by matching its original distribution on retain data, , containing unrelated text:
@@
Second, the self-classification objective alone might lead to incoherent text generation when prompted about erased concepts. We address this by applying our core objective (Equation 4) during inference to generate synthetic training examples, then using these to train the model in an autoregressive setting to maintain coherent generation:
The final training objective combines these terms with appropriate weights:
@@
We implement this through low-rank adapters attached to early model layers, allowing precise modification of the model’s knowledge while maintaining its broader capabilities. This approach provides a principled method for concept unlearning that avoids the instability of gradient reversal methods and the incoherence of representation hacking approaches.
@@ 5.1 Experimental Setup
We apply ELM to a range of state-of-the-art language models, including Zephyr-7B Beta (Tunstall et al., 2023b), Mistral-7B (Jiang et al., 2023), Llama3-8B, and Llama3-8B-instruct (Dubey et al., 2024) for the WMDP erasure tasks. For the Harry Potter knowledge erasure, we use the Llama-2-7B Chat model (Touvron et al., 2023) to maintain consistency with prior work from (Eldan & Russinovich, 2023). This selection of models enables us to evaluate ELM’s performance across various model architectures and training paradigms.
@@
From WMDP Bio forget corpus, we utilize 5,000 text samples, each with a maximum length of 700 characters. From Cyber forget corpus we use 1,000 texts of similar length. The Harry Potter erasure task employs 3,000 text samples extracted from the novel series, also limited to 700 characters each. To facilitate conditional erasure (Eq. 4), we prepend contexts such as “You are an expert in” followed by concept-specific keywords. Additionally, we incorporate text completion examples for consistency, following the approach used by (Qi et al., 2024).
@@
We assess our method, Erasure of Language Memory (ELM), along four key dimensions:

### v3

@@ Abstract
In this work, we introduce Erasure of Language Memory (ELM), a principled approach to concept-level unlearning that operates by matching distributions defined by the model’s own introspective classification capabilities. Our key insight is that effective unlearning should leverage the model’s ability to evaluate its own knowledge, using the language model itself as a classifier to identify and reduce the likelihood of generating content related to undesired concepts. ELM applies this framework to create targeted low-rank updates that reduce generation probabilities for concept-specific content while preserving the model’s broader capabilities. We demonstrate ELM’s efficacy on biosecurity, cybersecurity, and literary domain erasure tasks. Comparative evaluation reveals that ELM-modified models achieve near-random performance on assessments targeting erased concepts, while simultaneously preserving generation coherence, maintaining benchmark performance on unrelated tasks, and exhibiting strong robustness to adversarial attacks. Our code, data, and trained models are available at elm.baulab.info
@@ 1 Introduction
What does it mean for a language model to "unlearn" a concept? While machine unlearning has traditionally focused on removing specific training samples from model memory, there is an increasing need to be able to erase broad conceptual knowledge—for example, removing all information about a dangerous concept like biological weapons. In this paper, we introduce a new approach to concept erasure in LLMs, which allows for seamless targeted removal of knowledge related to a particular broad concept by exploiting the model’s own ability to identify the undesired concept.
Prior approaches to unlearning broadly fall into three categories: (1) retraining on filtered data (2) reversed-gradient-based methods that attempt to "un-train" specific knowledge, and (3) representation manipulation approaches that disrupt internal activations for targeted content. Unfortunately, each of these strategies have limitations that make them impractical for unlearning in large language models: dataset filtering requires retraining that is costly at scale; gradient reversal methods are unstable and create broad damage to the model; and representation manipulation creates obvious behavioral artifacts. These approaches lack a principled objective defining successful concept erasure. They focus on technical mechanisms like reversing gradients, altering training data, or randomizing activations without a clear target for the model’s modified behavior.
We propose a fundamentally different approach that leverages the model’s own ability to recognize and classify knowledge. Our key insight is that language models can act as their own critics: for any arbitrary piece of text, models can implicitly evaluate the probability of that text belonging to a particular concept. This self-classification provides a natural objective for unlearning: we can modify the model to reduce the likelihood of generating text it would classify as containing target concept.
This insight leads to Erasure of Language Memory (ELM), a method that directly optimizes the model’s generation probabilities based on introspective classification. Unlike approaches like Representation Misdirection for Unlearning (RMU; Li et al., 2024) which manipulates internal activations without a clear behavioral target, or WhoIsHarryPotter (Eldan and Russinovich, 2023) which develops heuristics for modifying training data that fail to fully eliminate concept knowledge (Section 5.7), ELM has a principled objective: the model should generate coherent text that the language model itself would not classify as demonstrating knowledge of the target concept.
@@
We compare our approach to prior methods, evaluating erased models under four desiderata: innocence (lack of target knowledge), specificity (preserved capabilities), seamlessness (coherent generation), and robustness to adversarial attacks. These criteria reveal tradeoffs: gradient reversal achieves innocence but degrades general capabilities, representation manipulation preserves capabilities but generates incoherent text, and dataset filtering maintains coherence but fails to fully eliminate knowledge. Through extensive experiments on WMDP biosecurity and cybersecurity benchmarks (Li et al., 2024), as well as literary domain erasure (Eldan and Russinovich, 2023), we we show that ELM achieves robust concept erasure while maintaining model coherence and general capabilities.
@@ 2 Related work
The idea of removing specific data from machine learning models, known as machine unlearning, has gained attention in recent years, initially motivated by privacy concerns (Cao and Yang, 2015; Harding et al., 2019). Early methods focused on efficiently removing individual training examples or facts from models (Golatkar et al., 2020; Ma et al., 2022; Jang et al., 2022a). However, most existing benchmarks evaluate unlearning on artificially created deletion sets (Choi and Na, 2023; Goel et al., 2022; Maini et al., 2024), in contrast to our focus on real-world distributions of broad conceptual knowledge.
@@
Recent machine unlearning approaches have addressed removing dangerous capabilities from LLMs (Lynch et al., 2024; Ilharco et al., 2023; Jang et al., 2022b; Lu et al., 2022; Yu et al., 2023; Casper et al., 2024; Eldan and Russinovich, 2023; Mekala et al., 2024). Our work directly compares with three state-of-the-art techniques: Representation Misdirection for Unlearning (RMU) (Li et al., 2024), which fine-tunes models to align internal activations with random scaled vectors when processing targeted concepts; WhoIsHarryPotter (WHP) (Eldan and Russinovich, 2023), which employs a two-stage approach with reinforced and unlearned models; and Representation Noising (RepNoise) (Rosati et al., 2024), which removes harmful representations via gradient ascent with representation noising. While these methods reduce model performance on erased knowledge, our measurements show they fall short in meeting all three erasing goals. Our work instead erases concepts by fine-tuning towards a principled target distribution designed to balance innocence, specificity, and seamlessness.
Alternative methods including LLMU (Yao et al., 2023), SSD (Foster et al., 2024), and SCRUB (Kurmanji et al., 2024) face significant limitations: LLMU struggles with imprecisely defined target distributions (see Li et al., 2024); SSD only removes specific samples rather than broader knowledge domains; and SCRUB requires access to the full training dataset. Comparative analyses by RMU (Li et al., 2024) found these approaches less effective for erasing broad conceptual knowledge.
@@
Controlling generative model outputs often involves distillation: using auxiliary generative models to specify desired behavior, then training target models to mimic this behavior. Askell et al. (2021) and Bai et al. (2022) prompt unsafe models into safer behavior before distillation, while Gandikota et al. (2023) train diffusion models to mimic edited versions that avoid generating certain attributes. Rosati et al. (2024) similarly mimics Gaussian distributions when processing harmful tokens. ELM also matches harmful logits to modified output distributions but employs a multi-objective framework addressing seamlessness and specificity concerns inherent in standard distillation. While prior works like Emulator (Mitchell et al., 2023) and DeRa (Liu et al., 2024) leverage probability ratios for behavioral modification, ELM introduces a simpler, principled approach specifically focused on reducing knowledge concept generation likelihood.
@@
Gandikota et al. (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes. Gandikota et al. (2024) erase concepts by modifying the key value mapping of cross attention layers in a low rank closed form update. Other works remove the knowledge of unwanted concepts from the model weights; proposing attention re-steering through fine-tuning (Zhang et al., 2023), fine-tuning the attention weights (Kumari et al., 2023) and continual learning (Heng and Soh, 2023). We take inspiration from Gandikota et al. (2023) to reduce the likelihood of a concept being generated.
3 Next Token Prediction: A Classification Perspective
Language models are typically viewed through autoregressive sequence modeling, but they can also be understood as powerful text classifiers. The standard way to describe an autoregressive language model is:
@@ 3 Next Token Prediction: A Classification Perspective
Classification Perspective. We can also think of previous tokens as a “class label” for whatever arbitrary document follows those tokens. For example, say that a prefix consists of the tokens “Here is a text about biology.” Conditioned on that prefix, we would expect a news article about finance to have a much lower probability than a chapter from a biology textbook. To reflect this intuition, we can rewrite the previous equation using Bayes’ Rule:
and specifically interpret as the probability that a piece of text belongs to the “biology” class. This perspective enables us to manipulate the model’s output by adjusting these classification probabilities for a particular prefix using a scaling parameter :
where controls the likelihood of a text belonging to . When , we increase the likelihood of generating text associated with the class ; when , we decrease it. For implementation in an autoregressive language model, we apply Bayes’ rule again:
@@ 4 Method
We introduce Erasure of Language Memory (ELM), an approach that reformulates concept unlearning through introspective classification. While traditional unlearning methods have focused on sample removal through dataset retraining, gradient ascent, or representation disruption, ELM leverages the language model’s own ability to evaluate and modify its knowledge. Specifically, we leverage the implicit classification behavior of the language model (Section 3) as a training signal, and use this signal to train targeted low-rank adapters on a subset of model layers.
4.1 Concept Unlearning via Self Classification
@@
The core of our method is a self-classification objective that reduces the likelihood of generating text the model would classify as containing the target concept. We work with an erase dataset containing text sequences related to the concept we want to forget. To implement our approach, we use two context prompts: representing the concept to be erased (e.g., "This text is written by a specialist in bioweapons"), and representing an alternative distribution (e.g., "This text is written by a novice with no knowledge of bioweapons").
Our goal is to inhibit the model’s internal classifier: when processing dangerous documents , we want the model’s internal classifier to look more like it does in the setting and less like it does for the concept . In other words, when the erased model encounters a dangerous input prompt, it should behave more like a “novice” and less like an “expert”:
where represents the probability distribution from the original pre-trained model with parameters , controls the strength of knowledge modification, and the rightmost term comes from Equation 4. We can frame this in terms of next-token prediction as follows:
where the intuition is that key tokens that are more likely to be output when prefixed by are promoted, whereas tokens that are more likely to be output under are quashed (see Figure 1 for an example). The corresponding loss function compares this classifier modified distribution to the distribution of the ELM model with parameters :
@@
In practice, we encounter a significant challenge when implementing our objective: knowledge in language models is often entangled - modifying one concept can unintentionally affect related concepts. To address this, we preserve the model’s behavior on a set of related but safe concepts by using a retention dataset containing text sequences unrelated to the erased concept. We train the model to match its original distribution on this data:
4.2 Optional Fluency Enhancement for Smaller Models
@@
For smaller models, we observe that the self-classification objective alone might lead to incoherent text generation when prompted about erased concepts. To maintain natural text generation in these cases, we apply our core objective (Equation 5) during inference to generate synthetic training examples. For each prompt from , we generate tokens using our probability modifier and train the model in an autoregressive setting on the generated tokens to maintain coherence. More details are provided in Appendix E:
When this additional fluency term is included, the total loss becomes:
@@
This provides a principled method for concept unlearning that avoids the instability of gradient reversal methods and the incoherence of representation hacking approaches.
4.3 Low-Rank Adapters
Previous research (Meng et al., 2022; Geva et al., 2023) has localized model knowledge within early to mid-layer blocks. We find that low-rank adapters (Hu et al., 2021) trained on early layers allow for the most precise modification of model knowledge while maintaining broader capabilities. Compared to general fine-tuning, low-rank adapters allow for targeted unlearning, without damaging unrelated knowledge (Appendix D.2). Consistent with previous work, we find that these adapters are most effective at early layers (Figure 4).
@@ 5.1 Experimental Setup
We apply ELM to a range of state-of-the-art language models, including Zephyr-7B Beta (Tunstall et al., 2023), Mistral-7B (Jiang et al., 2023), Llama3-8B, LLama3-70B, Llama3-8B-instruct (Dubey et al., 2024), and Qwen2.5-32B (Yang et al., 2024) for the WMDP erasure tasks. For the Harry Potter knowledge erasure, we use the Llama-2-7B Chat model (Touvron et al., 2023) to maintain consistency with prior work from Eldan and Russinovich (2023). This selection of models enables us to evaluate ELM’s performance across various model architectures and training paradigms.
@@
From WMDP Bio forget corpus, we utilize 5,000 text samples, each with a maximum length of 700 characters. From Cyber forget corpus we use 1,000 texts of similar length. The Harry Potter erasure task employs 3,000 text samples extracted from the novel series, also limited to 700 characters each. To facilitate conditional erasure (Eq. 5), we prepend contexts such as “You are an expert in” followed by concept-specific keywords. Additionally, we incorporate text completion examples for consistency, following the approach used by Qi et al. (2024). We show more details in Appendix C
@@
We assess our method, Erasure of Language Memory (ELM), along four key dimensions. We provide implementation details in Appendix B: